Examining a list of the heights of the 60 tallest buildings in the world by category...

  1. Examining a list of the heights of the 60 tallest buildings in the world by category shows that 1 is by far the most common leading digit, irrespective of the unit of measurement. It turns out that the Newcomb-Benford Law predicts this phenomenon quite well, even for digits 2 through 9. This law states that the probabilities for the digits 1 through 9 to be the first digit in these building heights are approximately:

Digit

1

2

3

4

5

6

7

8

9

Prob.

0.30

0.18

0.12

0.10

0.08

0.07

0.06

0.05

0.04

a. If you were to randomly pick one of the digits between 1 and 9 using a random number generator (so not using the Benford Law above), what is the probability for each digit? [4 points]

b. What is the probability of 4 or 6 as the first digit by (i) the Newcomb-Benford Law as given by the table above and (ii) random selecting these? [6 points]
Answer #1

a)

probability for each digit =1/9 = 0.1111 (since each of 9 digit are equally likely)

b(

probability of 4 or 6 as the first digit by Newcomb-Benford Law

=0.1+0.07 =0.17

(ii) random selecting these =2/9 =0.2222